OpenStudy (anonymous):

Solve for x: 2|x − 3| + 1 = 7

OpenStudy (anonymous):

Subtract 1 from both sides !

OpenStudy (anonymous):

Then multiply everything by 2 .

OpenStudy (anonymous):

i got x=6 but whats the second solution?

OpenStudy (callisto):

Not really... @Caramel_823 2|x − 3| + 1 = 7 1. Subtract 1 from both sides 2|x − 3| + 1 -1 = 7 - 1 2|x − 3| = 6 2. Divide both sides by 2 2|x − 3| /2= 6/2 |x − 3| = 3 3. remove the absolute sign x − 3 = 3 or x − 3 = -3 <- solve these 2 equations.

OpenStudy (anonymous):

thanks!

OpenStudy (callisto):

You're welcome! And welcome to OpenStudy too!